OpenStudy (anonymous):

Solve the linear system of equations 2x-3y=2 x=y-2

OpenStudy (anonymous):

Is ubstitution oka or am I thinking of the wrong thing?

OpenStudy (anonymous):

*substitution okay

OpenStudy (anonymous):

No its no substitution.

OpenStudy (anonymous):

Not sure what you mean then. Are you trying to find hat x and y equal?

OpenStudy (anonymous):

yes I'm trying to find x and y

OpenStudy (anonymous):

Okay, so what's wrong with... 2x-3y=2 x=y-2 2(y-2)-3y=2 2y-4-3y=2 -y=6 y=-6

OpenStudy (anonymous):

Then for x... x=y-2 x=-6-2 x=-8

OpenStudy (anonymous):

There are other methods she could use, I suppose. But I mean without her telling us what those methods are that she's using, we can't do much for her.

OpenStudy (anonymous):

True. I figured substitution would be the easiest one because we already had x=y-2. That and i couldn't remember the names of the others. :P